OpenStudy (anonymous):

Use sum and difference identities to find the exact value of sin 75 degrees. If someone could walk me through it that would be great. I have the equation but I don't know if im doing it right.

OpenStudy (anonymous):

sin 75 =sin(60 +15)=sin60cos15=cos60sin15 further sin15=sin(45-30) cos15=cos(45-30)

OpenStudy (amistre64):

what have you tried? that way we can tell if you are doing it right or not :)

OpenStudy (anonymous):

Um well im not really sure these pictures are all they have given me on how to do this.